OpenStudy (ptvrazz):

Which ordered pair is the solution to the system of equations? -3x+4y=-20 x+4y=-4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (serenity74):

Okay my friend do you know how to do system of equations?

OpenStudy (serenity74):

Cause none of us will give you the answer (welll most of us won't) so you have to tell me what you did so far so I can guide you through this awesome problem.

OpenStudy (ptvrazz):

i havent done anything i dont want the answer i need help to know how to do it.

OpenStudy (serenity74):

Okay good. Here is an example -8y + 4x = -10 2y +4x = 12 First choose which letter you wanna solve for doesn't matter which one you choose. I usually go with X, I get rid of X. -8y +4x = -10 2y + 4x= 12 in order to get rid of the X you have to multiply one equation by the negative sign since both of the equations are positive so it looks like this -8y+ 4x = -10 (-) [2y + 4x = 12] (distribute the negative sign) -8y + 4x = -10 -2y -4x =-12 then you get -10x = -22 x = -22/-10 = 22/10 = 11/5 then you plug the x value back into one of the orignial equations and solve for y 2y + 4(11/5) = 12 and then you get 2y = 12- 44/5 and solve for y

OpenStudy (serenity74):

Then you get the two functions. I chose a horrible example by the way but that's how you basically do it. If you need anymore help or clarifiation please tag me or let me know okay? Thank you :)

OpenStudy (serenity74):

I am soo sorry if i made it worse

OpenStudy (ptvrazz):

im sooooo confused gosh.

OpenStudy (serenity74):

Okay how I explain this to you hmmm wait I will give you this one website one sec

OpenStudy (serenity74):

-3x+4y=-20 x+4y=-4 right since you didn't specify the process I can do substitution method you make something equal to something so x = -4y - 4 then subsitutue that into the first equation's x value and solve for y -3(-4y -4) = -20 Y= ? distribute then solve for y after you are DONE with solving for Y you USE the Y-VALUE YOU GOT and plu it back into EITHER equation's Y-VAULE AND SOLVE for the X-VALUE X+ 4(?) = -4 NOW you solve for x here X =/ ( / , ? ) would be your answers (the numbers ofc)

OpenStudy (serenity74):

Any better or no???

OpenStudy (serenity74):

I feel really bad now :( I am so sorry